A laptop computer is purchased for $2600. Each year, its value is 80% of its value the year before. After how many years will the laptop computer be worth $900 or less?
Answer by Alan3354(69443)

A laptop computer is purchased for $2600. Each year, its value is 80% of its value the year before. After how many years will the laptop computer be worth $900 or less?
-----------------
After 1 year, it's $2600*0.8 = $2080
After 2 years, it's $2080*0.8 = $1664
etc
-------------
FV = $2600*(0.8^t) --- t in years
$900 = $2600*0.8^t
9/26 = 0.8^t
Log(9/26) = t*log(0.8)
t = log(9/26)/log(0.8) = ~4.7542 years
---> 5 years
=================
Not a well stated problem.
The answer can be "after 10 years" or after any # of years t > 4 (t as an integer).
You can answer "after 1000 years" and not be wrong.